When P = 100N and F = 85N, the box accelerates. The mass of the box is 25kg. Calculate the acceleration of the box.
1
Expert's answer
2020-04-27T10:04:39-0400


The net force acting on box


"F_{\rm net}=P-F=100-85=15\:\rm N."

The acceleration of the box


"a=\frac{F_{\rm net}}{m}=\frac{15}{25}=0.6\:\rm m/s^2."